40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 234 of 330  Not logged in ELOG logo
ID Date Author Type Category Subject
  4855   Wed Jun 22 15:24:10 2011 kiwamuUpdateABSLgot a laser controller for LightWave

Peter King came over to the 40m with a laser controller and gave it to us.

We will test it out with the LightWave NPRO, which was used for MOPA.

  4854   Wed Jun 22 12:29:57 2011 IshwitaSummaryAdaptive FilteringWeekly summary

I started on the 16th with a very intense lab tour & was fed with a large amount of data (I can't guarantee that I remember everything....)

Then... did some (not much) reading on filters since I'm dealing with seismic noise cancellation this summer with Jenne at the 40m lab.

I'll be using the Streckeisen STS-2 seismometers & I need to use the anti aliasing filter board that has the 4 pin lemo connectors with the seismometers & its boxes that require BNC connectors. I spent most of the time trying to solder the wires properly into the connectors. I was very slow in this as this is the first time I'm soldering anything.... & till now I've soldered 59 wires in the BNC connectors....

 

 

  4853   Wed Jun 22 12:24:44 2011 NicoleSummarySUSMidweek 2 Work Summary

I have made my transfer function model and posted it to the suspension wiki. Here is the link to my model!

Bode Plot Model

Please let me know if there need to be any adjustments, but I have posted the bode plots, a model image, and an explanation of why I think it's right! ^ ___^ V

I am currently working on the photo sensor circuit for the displacement detector. So far, I have gotten the infared LED to light up! ^ ___^ V

I am now trying to get a plot of forward voltage versus current for the LED. HOPEFULLY it will match the curve provided in the LED datasheet.

I'm using the bread board circuit box and when I'm not working at the bench, I have signs posted. PLEASE DO NOT REMOVE THE CONNECTIONS! It is

fine to move the bread board circuit box, but please do not disturb the connections > ____<

Here is a photo of the workspace

P6220200.JPG

  4852   Wed Jun 22 01:59:43 2011 SonaliUpdateGreen Lockingfibre-coupling of the IR beam

 What I did today.

1. Collimation of a beam.

  • I then practiced collimation of a 700 nm laser (output) beam after being coupled through a fibre.
  • I put together the set-up as shown in the attached picture where I used ....... to couple 650nm light into the PM.... fiber.
  • I kept shifting the focus of the output beam to an appreciable distance till it was approximately collimated.

2. Coupling of the IR light at the ETMY table to a fibre.

  • The fibre coupler was put in place to couple light into the fiber.
  • I put in the mirrors as planned to direct the IR beam exiting the doubling crystal towards the fiber coupler (input).
  • The mirrors were aligned such that the beam falls on the input lens of the coupler.
  • The far-end of the fibre originally would have gone to to PSL table but it has been put on this table to study the power of the IR beam transmitted through this set-up.  The output end of the fiber has been connected to another fiber optic coupler to collimate the exiting beam.
  • The picture of the current status attached.
  4851   Tue Jun 21 23:29:41 2011 kiwamuUpdateLSCsensing matrix measurement

I am now measuring the sensing matrix in the DRMI configuration.

A goal of tonight is to measure the sensing matrix as a test of the script.

 

The result will be updated later.

  4850   Tue Jun 21 20:35:50 2011 kiwamuUpdateLSCa script to measure sensing matrix

Last night I was making a script which will measure the sensing matrix using the realtime LOCKIN module.

The script is a kind of expansion of Jamie's one, which measure the asymmetry, to more generic purpose.

It will shake a suspended optic of interest and measure the response of each sensor by observing the demodulated I and Q signals from the LOCKIN module.

I will continue working on this.

 

  (current status)

 - made a function that drives the LOCKIN oscillator and get the data from the I and Q outputs.

 - checked the function with the MICH configuration.

   ITMX, ITMY and BS were shaken at 100 Hz and at different time.

   Then the response of AS55_Q showed agreement with what I got before for the actuator calibration (see this entry).

   It means the function is working fine.

  4849   Tue Jun 21 19:54:33 2011 SureshUpdateGreen LockingLightWave NPRO power supply shifted to ETMY end table

The Lightwave NPRO power supply which is being shared between the AS table and the ETMY table has been shifted back to the ETMY table

The current to the laser is set at 1.5A.  The laser output is 200mW at this current level.

  4848   Tue Jun 21 19:10:29 2011 SureshUpdateEnvironmentETMY end table

[Suresh, Sonali]

There were small pieces of glass, remnants of a fluorescent tube, which were lying around on the ETMY end table for a while now.  We picked up the larger pieces by hand and used the HEPA filtered vacuum cleaner to pick up the remaining glass and dust on the table. 

  4847   Tue Jun 21 16:11:13 2011 steveUpdateSAFETYsurf safety training

40m surfs:  Nicole Ing, Iswita Saikia and Sonali Mohapatra received 40m specific safety training  today.

  4846   Tue Jun 21 00:38:21 2011 SonaliUpdateGreen Lockingrepositioned "QPDY_PD"

1.The aim is the laser frequency stabilisation of PSL and AUX.

2.As a first step we want to couple some of the AUX laser beam into a single mode optical fibre and route the fibre to the PSL table.

3.The position of the optical fibre on the ETMY table is shown by the coupler in the attached picture. The yellow lines show the new scheme we want to implement.

4.WHAT WE DID TODAY.

  • The Y-arm was locked so that we could use the transmitted IR beam as the reference.
  • We shifted the position of the "QPDY_PD" .
  • We also shifted the "ETMYT" camera to make space for the "QPDY_PD".
  • The mirror  directing the beam into the "QPDY_PD" was rotated by 90 degrees to adhere to the new position of the "QPDY_PD".
  •  The attached photo shows the table as it is right now after the repositioning.
5.We continue with the positioning of the fibre-coupling tomorrow.
  4845   Mon Jun 20 18:36:49 2011 SureshUpdateLSCREFL55 PD update

[Suresh, Koji]

   I used a matlab code written by Koji to analyse the transimpedance and current noise data  of REFL55.  The details are in the attached pdf file.

Resonance is at 55.28 MHz:

Q of 4.5, Transimpedance of 615 Ohms

shot noise intercept current = 1.59 mA

current noise =21 pA/rtHz

 

Notch at 110.78 MHz:

Q of 54.8 Transimpedance of 14.68 Ohms.

 

 

 

 

 

 

 

 

 

Quote:

[Rana, Koji]

REFL55 was modified. The noise level confirmed. The PD is now ready to be installed.

 


Kevin's measurement report told us that something was wrong with REFL55 PD. The transimpedance looked OK, but the noise level was terrible (equivalent to the shotnoise of 14mA DC current).

Rana and I looked at the circuit, and cleaned up the circuit, by removing unnecessary 11MHz notch, 1k shunt resister, and so on.

I made a quick characterization of the PD.

First page:

The transimpedance ws measured as a function of the frequency. The resonance was tuned at 55MHz. The notch was tuned at 110MHz in order to reject the second harmonics. The transimpedance was ~540V/A at 55MHz. (For the calibration, I believed the DC transimpedance of 50V/A and 10000V/A for the DC paths of this PD and #1611, respectively, as well as the RF impedance (700V/A0 of #1611.

Second page:

Output noise levels were measured with various amount of photocurrent using white light from a light bulb. The measurement was perforemed well above the noise level of the measurement instruments.

Third page:

The measured output noise levels were converted into the equivalent current noise on the PD. The dark noise level agrees with the shot noise level of 1.5mA (i.e. 22pA/rtHz). In deed, the noise level went up x~1.5 when the photocurrent is ~1.4mA.

 

  4844   Mon Jun 20 18:12:20 2011 NicoleUpdateSUSSmall Table Cleaned and Levelled

P6220198.JPG

The small optical bench (next to the MC-2 Chamber and the tool box tower) has been cleared of the misc. object previously on it, cleaned, and leveled (after much calibration X___X).

PLEASE, PLEASE, PLEASE do NOT MOVE OR HIT THE TABLE! It was incredibly painful to level.

This is how leveling the table made me feel...

P6220199.JPG

VERY SAD...so do not move please!

The shaker has already been moved to the table and the amplifier for my shaking experiment is located behind the table (not on the table, as to prevent scratching).

 

 

  4843   Mon Jun 20 17:58:00 2011 ranaUpdateCDSGateway program killed

There was a rogue, undocumented, gateway process running on NODUS since ~4 PM. This guy was broadcasting channels back into the Martian and causing lockups in the IOO controls. I did a kill -9 on its process.

Someone will pay for this.

  4842   Mon Jun 20 16:44:02 2011 steveUpdateABSLI-P curve of LightWave M126-1064-700

 

 Put the serial numbers into the elog. So we can identify the laser and controller in the future.

The old days the NPRO ( inside the MOPA ) was running ~1.7A  500 mW

  4841   Mon Jun 20 13:48:25 2011 KojiUpdateABSLI-P curve of LightWave M126-1064-700

Hmm. Was the current within the operating range? (i.e. Is it a 700mW laser or a 1W one?)

You can obtain the nominal operating current from the old EPICS values (or some elog entries).

Note that NPROs are designed to be healthy only at around the nominal pumping power
(i.e. thermal gradient, and thermal lensing of the crystal, etc.)

ALSO:

Be aware that this laser should be used under the old SOP. So the appropriate interlocking is mandatory.

And probably we need to modify the SOP such that it reflects the latest situation.

Quote:

The I-P curve of the LightWave NPRO, which was taken out from the MOPA box, was measured. It looks healthy.

The output power can go up to about 1 W, but I guess we don't want it to run at a high power to avoid any further degradation since the laser is old.

  4840   Mon Jun 20 11:38:49 2011 kiwamuUpdateABSLI-P curve of LightWave M126-1064-700

The I-P curve of the LightWave NPRO (M126-1064-700), which was taken out from the MOPA box, was measured. It looks healthy.

The output power can go up to about 1 W, but I guess we don't want it to run at a high power to avoid any further degradation since the laser is old.

 

IP_curve.png

 X-axis is the current read from the display of the controller. Y-axis is the output power, directly measured by Coherent PM10.

The measurement was done by changing the current from the controller.

Quote from #4832

 [Things to be done]

   - measure the beam profiles and power

   - get a laser controller, which will be dedicated for this laser, from Peter King

  4839   Mon Jun 20 11:04:03 2011 NicoleUpdateSUSWork Plan for Week 2

Here is my work plan for this week:

Current Week Plan (Week 2) (As of 6/17/11)

 

Setting Up for Horizontal Displacement Measurements

1) Help Steve clean small table for experiment

2) Remove aluminum base from TT suspension

3) Mount shaker onto table base

4) Mount horizontal slider onto table base

5) Connect TT suspension, shaker, and horizontal slider

Begin Assembly of Sensors

1) Begin building circuit for displacement photosensors

2) Calibrate photosensor using linear regions of power versus distance curves

3) Circuit box for photosensors?

  4838   Mon Jun 20 10:45:43 2011 JamieUpdateCDSPower restored to 1X1/1X2 racks. IOO binary output module installed.

All power has been restored to the 1X1 and 1X2 racks.  The modecleaner is locked again.

I have also hooked up the binary output module in 1X2, which was never actually powered.  This controls the whitening filters for MC WFS.  Still needs to be tested.

  4837   Mon Jun 20 09:28:19 2011 JamieUpdateCDSShutting down low-voltage DC power in 1X1/1X2 racks

In order to install the BO module in 1X2, I need to shut down all DC power to the 1X1 and 1X2 racks.

  4836   Mon Jun 20 09:04:13 2011 steveUpdateSUSETMX damping restored

ETMX sus damping restored

  4835   Mon Jun 20 00:59:02 2011 kiwamuSummaryGeneralWeekly report
This is a summary for the week ending June 19th. Feel free to edit this entry.
(Number of elog entries = 27)

* Refinement of LSC screen
    -> Kissel buttons and some indicators were newly installed
    -> A script to autonatically generate kissel buttons was made

* New BIO installed on ETMY

* LightWave for ABSL
    -> taken out from the MOPA box and put on the AP table with temporary use of the Y end laser controller
 
* Shipping 2 RFPDs to LLO
 
* LEDs on the BIO for the vertex suspensions were blown
    -> fixed and re-installed. A test script will be prepared
 
* PEM AA board was fixed
 
* A plot of the MICH noise was produced for the first time
 
* Schnupp asymmetry measurement : Las = 3.64+/-0.32cm
 
* The photo diode on WFS2 has been replaced by YAG-444-4A
 
* SUS binary IO crates were taken out
 
* Fiber died
     ->C1LSC was unable to communicate to C1SUS. Installing a new copper Dolphine fixed the issue.
 
* SURF students came
  4834   Fri Jun 17 23:20:05 2011 KojiUpdateLSCSome updates of the LSC screen

Some updates of the LSC screen

- Signal amplitude monitor for the PD signals (--> glows red for more than 1000)

- Kissel Buttons for the main matrices

- Trigger display at the output of the DOF filters

- Signal amplitude monitor for the SUS LSC output (--> glows red for more than 10000)

 

ADC Over flow monitor is showing some unknown numbers (as ADCs are handled by IOPs).
I asked Joe for the investigation (and consideration for the policies)

  4833   Fri Jun 17 17:02:15 2011 JamieUpdateSUSETMX/ETMY binary output modules (re)installed, not yet tested

I have installed a new binary output module in ETMY, where there was none previously.  It is installed, powered (with working LEDs), hooked up (to the binary output card and the cross connect), but it hasn't been fully tested yet.

I also re-installed the binary output module in ETMX, with newly modified power-indicator LEDs.

Both modules are fully installed, but they have not yet been fully tested to confirm that they are indeed switching the whitening and de-whitening filters.

  4832   Fri Jun 17 16:05:07 2011 kiwamuUpdateABSLLightWave out of MOPA box

[Suresh / Kiwamu]

 We did the following things :

   - Took the LightWave NPRO out from the MOPA box

   - Temporarily took out the laser controller which has been connected to the Y end laser.

   - Put the LightWave on AP table and plugged the laser controller and confirmed that it still emits a beam

 DSC_3139_small.png

 

[Things to be done]

   - measure the beam profiles and power

   - get a laser controller, which will be dedicated for this laser, from Peter King

 

[Background and Motivation]

 The PRC and SRC length have to be precisely measured before the vent.

In order to measure those absolute length we are going to use the Stochino technique, which requires another laser to scan the cavity profiles.

The LightWave NPRO laser in the MOPA box was chosen for the Stochino laser because it has a large PZT range of 5 MHz/V and hence allows us to measure a wider frequency range.

The laser in the MOPA box had been connected to home-made circuits, which are not handy to play with. So we decided to use the laser with the usual laser controller.

Peter King said he has a LightWave laser controller and he can hand it to us.

Until we get the controller from him we do some preparations with temporary use of the Y end laser controller.

  4831   Fri Jun 17 08:03:48 2011 steveUpdateSUSITMY sus damping restored

ITMY sus damping restored.

 

  4830   Fri Jun 17 00:17:26 2011 ranaConfigurationElectronics2 RFPDs sent to LLO

Koji and I found 2 RFPD boxes to send to LLO. We've put them onto Steve's desk to be overnighted to Valera.

One of them is our old 21.5 MHz gold box RFPD from the FSS (which we don't use). The other one is a 2mm gold box one which was previously tuned for 66 MHz.

 

They shipped out on Friday

  4829   Thu Jun 16 23:19:09 2011 KojiUpdateSUSVertex SUS Binary Output Boxes removed

[Jamie, Koji]

- We found the reason why some of the LEDs had no light. It was because the LEDs were blown as they were directly connected to the power supply.
The LEDs are presumably designed to be connected to a 5V supply (with internal current-limiting resistor of ~500Ohm). The too much current
with the 15V (~30mA) made the LED blown, or the life-time of them shorter.

- Jamie removed all of the BO modules and I put 800Ohm additional resister such that the resultant current is to be 12mA.
The LEDs were tested and are fine now.

- The four BO boxes for C1SUS were restored on the rack. I personally got confused what should be connected where
even though I had labeled for BO0 and BO1. I just have connected CH1-16 for BO0. The power supplies have been connected only to BO0 and BO1.

- I tested the whitening of PRM UL sensor by exciting PRM UL sensor. The transfer function told us that the pendulum response can be seen
up to 10-15Hz. When the whitening is on, I could see the change of the transfer function in that freq band. This is good.
So the main reason why I could not see theis was that the power supply for the BOs were not turned on.

- I suppose Jamie/Joe will restore all of the BO boxes on the racks tomorrow. I am going to make a test script for checking the PD whitenings.

  4828   Thu Jun 16 08:45:14 2011 steveUpdateSUSVertex SUS Binary Output Boxes removed

Quote:

- I was investigating the SUS whitening issue.

- I could not find any suspension which can handle the input whitening switch correctly.

- I went to 1X5 rack and found that both of the two binary output boxes were turned off.
As far as I know they are pulling up the lines which are switched by the open collector outputs.

- I tried to turn on the switch. Immediately I noticed the power lamps did not work. So I need an isolated setup to investigate the situation.

- The cables are labelled. I will ask steve to remove the boxes from the rack.

 I shut down damping to the Vertex optics and removed Binary IO  Adapter chassy BO0 and BO1

About a week ago I discussed the BO0's power indicator lights with Kiwamu. They were  not on or they were blinking on-off.

I put screws into ps connectors in the back, but it did not helped.

  4827   Thu Jun 16 00:43:36 2011 KojiUpdateSUSVertex SUS Binary Output Boxes were turned off / need investigation

- I was investigating the SUS whitening issue.

- I could not find any suspension which can handle the input whitening switch correctly.

- I went to 1X5 rack and found that both of the two binary output boxes were turned off.
As far as I know they are pulling up the lines which are switched by the open collector outputs.

- I tried to turn on the switch. Immediately I noticed the power lamps did not work. So I need an isolated setup to investigate the situation.

- The cables are labelled. I will ask steve to remove the boxes from the rack.

  4826   Thu Jun 16 00:39:08 2011 KojiUpdatePEMPEM AA Board has been diagnosed and fixed

As seen in the photo, the board has a strange bulge on the board,
and
the color of the internal line around the bulge got darkened.

I don't trust this board any more. We should switch to the alternative one.

Quote:

Steve will re-install the board in the rack in the morning.

 

  4825   Wed Jun 15 16:37:56 2011 JenneUpdatePEMPEM AA Board has been diagnosed and fixed

[Jenne, Steve]

After talking with Steve, I had a look at the PEM's AA board, to see what the problem was. 

Steve said the symptom he had noticed was that the Kepco power supplies which supply the +\- 5 V to the AA board were railing at their current limits as soon as he plugged the board in.  Also, he smelled smoke. 

I started with the power supplies, and saw that the 2 individual supplies each had a dV=5V, and that the one labeled +5V had the red wire on the + output of the power supply, and the black wire on the - output.  The supply labeled -5V had the orange wire on the -output of the power suppy, and the black wire on the + output.  Normally, you would expect that the 2 black wires are also connected together, and perhaps also to ground.  But at least together, so that they share a common voltage, and you get +\- 5V.  However these 2 power supplies are not connected together at all. 

This implies that the connection must be made on the AA boards, which I found to be true.  It seems a little weird to me to have that common ground set at the board, and not at the power supplies, but whatever.  That's how it is.

The problem I found is this:  The keyed connectors were made backward, so that if you put them in "correctly" according to the key, you end up shorting the +5V to the -5V, and the 2 black wires are not connected together.  You have to put the keyed connectors in *backwards* in order to get the correct wires to the correct pins on the board.  See the attached pdf figure.

Since these are internal board connections, and they should not ever be changed now that Steve has put in the adapter thing for the SCSI cable, I'm just leaving them as-is.  Steve is going to write in huge letters in sharpie on the board how they're meant to be connected, although since this problem wasn't caught for many many years, maybe it won't ever be an issue again.  Also, we're going to move over to the new Cymac system soon-ish.  However, whomever made the power cable connector from the box to the board for this AA board was lazy and dumb.

After putting the connectors on the way they needed to be, Steve and I powered up the board, hooked up the SCSI cable in the back, and put a constant voltage (~1.3VDC battery) across various different channels, and confirmed that we could see this voltage offset in Dataviewer. (Kiwamu is hoarding both of our SRS function generators, so we couldn't put in a low freq sine wave like I normally would). Everything looked okie dokie, so I'll check the regular PEM channels tomorrow.

Steve will re-install the board in the rack in the morning.

  4824   Wed Jun 15 15:18:01 2011 kiwamuUpdateGeneralWednesday cleaning

[Jenne / Kiwamu]

We spent approximately an hour for the weekly Wednesday cleaning.

This time we moved onto an area where a desk and optics shelf reside along the Y arm.

We will continue cleaning up there in the next time too.

  4823   Wed Jun 15 12:16:37 2011 JamieUpdateLockingMICH noise budget?

Quote
If you tilt your head sideways, you will notice that in this plot (totally uncalibrated, as yet), the BLACK trace, which is my white-light measurement of the AS55 shot noise is above the AS55Q noise when the Michelson is locked (true only at low frequency).  You will also notice that the same appears to be true for the Whitening Filter + Antialiasing Filter + ADC noise (GRAY trace).  Since Black, Gray, Pink and Green should all have the same calibration factor (a constant), calibrating the plot will not change this.  Brown and Blue are the MICH_OUT (aka MICH_CTRL) for dark and bright fringes, respectively.

Hey, Jenne.  I think there are a couple of things.  First, you're missing a PD dark noise measurement, which would be useful to see.

But I think the main issue is that it sounds like all of your closed loop measurements are done with the in-loop PD.  This means that everything will be suppressed by the loop gain, which will make things look like they have a noise lower than the actual noise floor.

  4822   Wed Jun 15 02:20:00 2011 JenneUpdateLockingMICH noise budget?

I would like to announce my confusion with regard to the MICH noise budget, in hopes that someone else has some inspiration

If you tilt your head sideways, you will notice that in this plot (totally uncalibrated, as yet), the BLACK trace, which is my white-light measurement of the AS55 shot noise is above the AS55Q noise when the Michelson is locked (true only at low frequency).  You will also notice that the same appears to be true for the Whitening Filter + Antialiasing Filter + ADC noise (GRAY trace).  Since Black, Gray, Pink and Green should all have the same calibration factor (a constant), calibrating the plot will not change this.  Brown and Blue are the MICH_OUT (aka MICH_CTRL) for dark and bright fringes, respectively.

I measure 58mV at the DC out of the AS55 PD when the Michelson is locked on the bright fringe.  This (assuming DC transimpedance of 50ohms) gives 1.16mA of DC photo current.

So.  What is going on here?  Am I totally confused??

In other news, assuming (which I'm not 100% confident about right now) that these traces are vaguely correct, the Michelson is limited by shot noise above ~20Hz.  This is...good?  We want to be shot noise limited.  Do we want to be limited at such a low frequency?

(Also, yes I can calibrate the plot to m/rtHz, but no, I won't tonight because something is funny with my calibration for the free running noise and I'll fix it tomorrow.)

MICH_noise_budget_measurements_15June2011.jpg

  4821   Wed Jun 15 01:30:38 2011 JamieSummaryLSCSchnupp asymmetry measurement

Measurement of Schnupp asymmetry

This was done by measuring the relative phase between the sidebands reflected from the two arms while the arm cavities are locked.

The Schnupp asymmetry is measured to be:   Lsa = 3.64 ± 0.32 cm

schnupp.png

Description:

As a phase reference we use the zero crossing of the response function for the out-of-phase control signal for the single arm cavity lock [0]. The difference in the RD rotation phase of the response zero crossings indicates the phase difference in the sideband signals reflected from the arms. Assuming the asymmetry is less than half the RF modulation wavelength [1], the asymmetry is given by the following formula:

       \Delta \phi   c   1 
L_sa = ----------- ----- -
           360     f_RSB 2

We use a LSC digital lock-in to measure the response of the arm cavity at a single-frequency drive of it's end mirror.

[0] The locations of the zero crossings in the out-of-phase components of the response can be determined to higher precision than the maxima of the in-phase components.

[1] fRSB = 55 MHz,     c/fRSB/2 = 2.725 m

Procedure:

  1. Lock/tune the Y arm only.
    • We use AS55_I to lock the arms.
  2. Engage the LSC lock-in.
  3. Tune the lock-in parameters:
  4. lock-in freq: 103.1313 Hz
    I/Q filters:  0.1 Hz low-pass
    phase:        0 degrees
    
  5. Set as input to the lock-in the out-of-phase quadrature from the control RFPD.  In this case AS55_Q->LOCKIN.
  6. Drive the arm cavity end mirror by setting the LOCKIN->Y_arm element in the control matrix.
  7. Note the "RD Rotation" phase between the demodulated signals from the control PD (AS55)
  8. For some reasonable distribution of phases around the nominal "RD Rotation" value, measure the amplitude of the lock-in I output.
    • Assuming the Q output is nearly zero, it can be neglected.  In this case the Q amplitude was more than a factor of 10 less than the I amplitude.
    • Here we take 5 measurements, each separated by one over the measurement bandwidth (as determined by the lock-in low pass filter), in this case 10 seconds.  The figure above plots the mean of these measurements, and the error bars indicate the standard deviation.

The data and python data-taking and plotting scripts are attached.

Error Analysis:

To to determine the parameters of the response (which we know to be linear) we use a weighted linear least-squares fit to the data:

y = b X

where:

X0j = 1
X1j = xj              # the measurement points
y = yi                 # the response
b = (b0, b1)     # line parameters

The weighting is given by the inverse of the measurement covariance matrix. Since we assume the measurements are independent, the matrix is diagonal and Wii = 1/\sigmai2 The
estimated parameter values are given by:

\beta  =  ( XT W X )-1 XT W y  =  ( X'T X' )-1 X'T y'

where X' = w X, y' = w y and wii = \sqrt{Wii}.

The X' and y' are calculated from the data and passed into the lstsq routine. The output is \beta.

The error on the parameters is described by the covariance matrix M\beta:

M\beta = ( XT W X)-1 = ( X'T X')-1

with correlation coefficients \rhoij = M\betaij / \sigmai / \sigmaj.

The x-axis crossing is then given by:

X(Y=0) = - \beta1 / \beta0

References:

Valera's LLO measurement
http://en.wikipedia.org/wiki/Weighted_least_squares
http://en.wikipedia.org/wiki/Linear_least_squares_(mathematics)#Weighted_linear_least_squares
http://en.wikipedia.org/wiki/Error_propagation

  4820   Wed Jun 15 00:50:11 2011 KojiHowToComputer Scripts / ProgramsKissel Button Generator

Now the Kissel-button generator takes the command line arguments and options.
The script is fully documented by the usage message of the script itself.
It still needs the external supporting files "MATRIX*.adl_parts".

Now the LSC screen has these buttons for the input and output matrices.
The command lines to generate those buttons are listed at the end of this entry as the examples.


>pwd
/opt/rtcds/caltech/c1/medm/c1lsc_tst/master/KisselButtonGenerator

>./generate_KisselButton.py -h
usage:
generate_KisselButton.py [options]  end_row end_column matrix_ch_name

This generates an MEDM screen of a button with the style designed by
Jeff Kissel for his ISI screens. This button has a display of a matrix
elements. If the matrix element is non-zero it glows in green. Otherwise
its color is dark. Usually the button created by this script
is to be copy-pasted to other screens.

Three arguments have to be given:
  end_row         the number of the row at the end
  end_column      the number of the column at the end
  matrix_ch_name  the channel name of the matrix to be monitored
                  e.g. give C1:LSC-OUTPUT_MTRX for C1:LSC-OUTPUT_MTRX_1_1, ...

There are options prepared in order to control the parameters of the button.

example:
generate_KisselButton.py 6 6 C1:LSC-OUTPUT_MTRX
      6x6 matrix for C1:LSC-OUTPUT_MTRX


options:
  -h, --help          show this help message and exit
  --sr=START_ROW      specify the starting row number for the button array.
                      [default: 1]
  --sc=START_COLUMN   specify the starting column number for the button array.
                      [default: 1]
  --bw=BUTTON_WIDTH   specify the pixel width of the small button. [default:
                      8]
  --bh=BUTTON_HEIGHT  specify the pixel height of the small button. [default:
                      8]
  --dl=DISPLAY_LABEL  specify the button label. [default: channel name]
  --sn=SCREEN_NAME    specify the file name of the screen opened when one
                      click the button. The relative or absolute path can be
                      included. [default: a name guessed from the channel
                      name. e.g. C1LSC_OUTPUT_MTRX.adl for C1:LSC-OUTPUT_MTRX]

>./generate_KisselButton.py --bw=3 --bh=4 --dl="RFPD InMTRX" 16 8 C1:LSC-PD_DOF_MTRX > rfpd_mtrx.adl

>./generate_KisselButton.py --sc=21 --bw=6 --bh=4 --dl="DCPD InMTRX" 27 8 C1:LSC-PD_DOF_MTRX > dcpd_mtrx.adl

>./generate_KisselButton.py --bw=4 --bh=4 --dl="Trig MTRX" 11 8 C1:LSC-TRIG_MTRX > trig_mtrx.adl

>./generate_KisselButton.py --bw=4 --bh=4 --dl="Out MTRX" 9 10 C1:LSC-OUTPUT_MTRX > output_mtrx.adl

  4819   Wed Jun 15 00:49:34 2011 SureshUpdateIOOWFS2 has been fixed.

 

The WFS2 sensor head had a damaged Quadrant PIN diode (YAG-444-4A).  This has been replaced by a   YAG-444-4AH  which has a responsivity of 0.5 A/W. 

P6150121.JPG     P6150124.JPG

The responsivity of each quadrant was measured at normal incidence.  A diagram of the set up with the relevant power levels is attached.  The precision of these measurement is about 5% .  Largely because the power levels measured are sensitive to the position of the laser beam on the power meter sensor head (Ophir with ND filter mask taken off).  Putting the mask back on did not solve this problem.

The incident power was 0.491mW  of which about 0.026mW was reflected from the face of the QPD.  The beam was repositioned on the QPD to measure the response of each quadrant.  In each case the beam was positioned to obtain maximum DC output voltage from the relevant quadrant.  A small amount of spill over was seen in the other quadrants.  The measurements are given below

WFS2 DC output measurements (mV)
  Position 1 Position 2 Position 3 Position 4 Dark
Q1 244 6.7 5.4 6.9 4
Q2 5.9 238 8.4 5 5
Q3 9 6.6 236 7.3 6
Q4 7.5 7 7 252 7

WFS_QE_measurement.png

To measure these DC outputs of from the sensor-head a breakout board for the 25-pin D-type connector was used as in the previous measurements.  The results are given below

 

WFS2 Quantum Efficiency measurement

  DC out (mV)

Responsivity

A/W

Quantum Efficiency (%)
Q1 238 0.52 0.60
Q2 233 0.50 0.59
Q3 230 0.50 0.58
Q4 244 0.53 0.61

 

The measured responsivity agrees with the specification from the manufacturer.  It is to be noted that the previous QPD is reported to have a slightly smaller responsivity 0.4 A/W at 1064 nm.  The data sheet is attached. 

Since the new QPD may have a slightly different capacitance the RF transfer function of the WFS2 needs to be examined to verify the location of the resonances. 

 

Quote:

[Larisa and Jenne]

A few weeks ago (on the 28th of January) I had tried to measure the quantum efficiency of one quadrant of the WFS as a function of angle.  However, Rana pointed out that I was a spaz, and had forgotten to put a lens in front of the laser.  Why I forgot when doing the measurement as a function of angle, but I had remembered while doing it at normal incidence for all of the quadrants, who knows?

Anyhow, Larisa measured the quantum efficiency today.  She used WFS2, quadrant 1 (totally oil-free), since that was easier than WFS1.  She also used the Jenne Laser (with a lens), since it's more stable and less crappy than the CrystaLasers.  We put a 50 Ohm terminator on the RF input of the Jenne Laser, since we weren't doing a swept sine measurement.  Again, the Ophir power meter was used to measure the power incident on the diode, and the reflected power, and the difference between them was used as the power absorbed by the diode for the quantum efficiency measurement.  A voltmeter was used to measure the output of the diode, and then converted to current as in the quote below. 

Still on the to-do list:  Replace the WFS2 diode.  See if we have one around, otherwise order one.  Align beams onto WFS so we can turn on the servo.

QE = (h*c)/(lambda*e) * (I/P)

Where I = (Volts from Pin1 to GND)/2 /500ohms
P = Power from laser - power reflected from diode.
h, c, e are the natural constants, and lambda is 1064nm.
Also, I/P = Responsivity


Larissa is going to put her data and plots into the elog shortly....

Quote:

Quantum Efficiency Measurement:

I refer to Jamie's LHO elog for the equation governing quantum efficiency of photodiodes: LHO 2 Sept 2009

The information I gathered for each quadrant of each WFS was: [1] Power of light incident on PD (measured with the Ophir power meter), [2] Power of light reflected off the PD (since this light doesn't get absorbed, it's not part of the QE), and [3] the photo current output by the PD (To get this, I measured the voltage out of the DC path that is meant to go to EPICS, and backed out what the current is, based on the schematic, attached). 

I found a nifty 25 pin Dsub breakout board, that you can put in like a cable extension, and you can use clip doodles to look at any of the pins on the cable.  Since this was a PD activity, and I didn't want to die from the 100V bias, I covered all of the pins I wasn't going to use with electrical tape.  After turning down the 100V Kepco that supplies the WFS bias, I stuck the breakout board in the WFS.  Since I was able to measure the voltage at the output of the DC path, if you look at the schematic, I needed to divide this by 2 (to undo the 2nd op amp's gain of 2), and then convert to current using the 499 Ohm resistor, R66 in the 1st DC path.  

I did all 4 quadrants of WFS1 using a 532nm laser pointer, just to make sure that I had my measurement procedure under control, since silicon PDs are nice and sensitive to green.  I got an average QE of ~65% for green, which is not too far off the spec of 70% that Suresh found.

I then did all 8 WFS quadrants using the 1064nm CrystaLaser #2, and got an average QE of ~62% for 1064 (58% if I exclude 2 of the quadrants....see below).  Statistics, and whatever else is needed can wait for tomorrow.

Problem with 2 quadrants of WFS2?

While doing all of this, I noticed that quadrants 3 and 4 of WFS2 seem to be different than all the rest.  You can see this on the MEDM screens in that all 6 other quadrants, when there is no light, read about -0.2, whereas the 2 funny quadrants read positive values.  This might be okay, because they both respond to light, in some kind of proportion to the amount of light on them.  I ended up getting QE of ~72% for both of these quadrants, which doesn't make a whole lot of sense since the spec for green is 70%, and silicon is supposed to be less good for infrared than green.  Anyhow, we'll have to meditate on this.  We should also see if we have a trend, to check how long they have been funny.

 

 

  4818   Tue Jun 14 18:12:34 2011 Jamie, KiwamuUpdateLSCLSC seems to be fully recovered

We are now locking the arms reliably, with reasonable transmitted power.  We zeroed the LSC offsets with script, since they were apparently not being reset with either the overall burt restore or the arm restore scripts.

We have lost a bit of power through the mode cleaner.  However, we have opted not to tweak it up just yet, so that we don't have to realign to the arms.

  4817   Tue Jun 14 16:38:31 2011 JamieUpdatePSLTweaked input pointing to PMC

The PMC trans power was a little low (0.77V or so).  I tweaked up the input pointing and now we're getting about 0.875V transmitted.

  4816   Tue Jun 14 12:23:44 2011 Jamie, JoeUpdateCDSWE ARE ALL GREEN! LSC back up and running in new configuration.

After moving the c1lsc computer to 1X4, then connecting c1lsc to it's IO chassis in 1Y3 by a fiber PCIe extension cable, everything is back up and running and the status screen is all green.  c1lsc is now directly connected to c1sus via a short copper Dolphin cable.

After lunch we will do some more extensive testing of the system to make sure everything is working as expected.

  4815   Tue Jun 14 09:25:17 2011 JamieUpdateCDSDolphin fiber tested with c1sus, still bad

The bad Dolphin was still bad when tested with a connection between c1sus and the Dolphin swtich.

I'm headed over to Downs to see if we can resolve the issue with the PCIe extension fiber.

  4814   Tue Jun 14 09:24:36 2011 steveConfigurationPhotosSUS binary IO chassis 2 and 3 moved from 1X5 to 1X4

Quote:

While preping 1X4 for installation of c1lsc, we removed some old VME crates that were no longer in use.  This freed up lots of space in 1X4.  We then moved the SUS binary IO chassis 2 and 3, which plug into the 1X4 cross-connect, from 1X5 into the newly freed space in 1X4.  This makes the cable run from these modules to the cross connect much cleaner.

 Are we keeping these?

  4813   Tue Jun 14 03:15:29 2011 KojiHowToComputer Scripts / ProgramsKissel Button Generator

I have made a python script to generate the button designed by Jeff Kissel for his ISI screen.

It is currently located at the following location:
/cvs/cds/rtcds/caltech/c1/medm/c1lsc_tst/master/KisselButtonGenerator/generate_KisselButton.py
but should be relocated to somewhere appropriate.
It also uses fragmented medm files named "MATRIX*.adl_parts".

# Jamie, could you suggest the right place?

The parameters are assigned at the beggining of the script.
This script print the result to stdout. So you need to redirect the output into a file.
e.g.

> ./generate_KisselButton.py >tmp.adl

The script should be modified such that it accepts the command line options.
It needs more python learning for me.


# Number of the column
mat_h = 20;

# Number of the row
mat_v = 10;

# horizontal pixel size of the rectangular display for each matrix element
button_width = 8;

# vertical pixel size of the rectangular display for each matrix element
button_height = 8;

replace_dict = {
# Title
    '${DISPLAY_LABEL}':'ITMX_INMATRIX', 
# Path of the MEDM file to be open by clicking the button
    '${DISPLAY_NAME}':'/cvs/cds/rtcds/caltech/c1/medm/c1sus/master/C1SUS_ITMX_INMATRI
X_MASTER.adl',
# The channel name of the matrix element
# ($V and $H are replaced to the numbers i.e. "_3_4")
    '${MATRIX_CHAN}':'C1:SUS-ITMX_INMATRIX_$V_$H'
    };


 

  4812   Mon Jun 13 19:26:42 2011 Jamie, JoeConfigurationCDSSUS binary IO chassis 2 and 3 moved from 1X5 to 1X4

While preping 1X4 for installation of c1lsc, we removed some old VME crates that were no longer in use.  This freed up lots of space in 1X4.  We then moved the SUS binary IO chassis 2 and 3, which plug into the 1X4 cross-connect, from 1X5 into the newly freed space in 1X4.  This makes the cable run from these modules to the cross connect much cleaner.

  4811   Mon Jun 13 18:40:08 2011 Jamie, JoeUpdateCDSSnags in the repair of LSC CDS

We've run into a problem with our attempts to get the LCS control back up and running.

As reported previously, the Dolphin fiber connection between c1lsc and c1sus appears to be dead.  Since we have no replacement fiber, the solution was to move the c1lsc machine in to the 1X4 rack, which would allow us to use one of the many available short Dolphin cables, and then use a long fiber PCIe extension cable to connect c1lsc to it's IO chassis.  However, the long PCIe extension cable we got from Downs does not appear to be working with our setup.  We tested the cable with c1sus, and it does not seem to work.

We've run out of options today.  Tomorrow we're going to head back to Downs to see if we can find a cable that at least works with the test-stand setup they have there.

  4810   Mon Jun 13 16:27:10 2011 JenneUpdateElectronicsSeismometer Box Update/graphs

Quote from elog 4807:

The noise graphs relating total noise of the Seismometer circuit (GURALP stuff) to the LIGO seismic noise curve have been completed started.

 What Larisa meant to post (I'm sure) is something more like this (sorry it's a little squished...I put too many words in the legend):

I've only included the 2 noise contributions from the LISO model that seem to dominate the sum noise.  The plot gets a little crazy if you include all of the non-important sources.

NewSeisBoxNoise.png

So, what's the point??

First, the new box design doesn't have any crazy-special op-amps in it, so the noise of the new box is probably comparable to the old box.  So, if that's true, the old box may not have been limiting the differential seismic noise.  This definitely needs to be checked out.  I'll make a quickie LISO model of the old Guralp breakout box, to see what its noise actually looks like, according to LISO.  If it wasn't ever the breakout box that was limiting us, what the heck was it??

Second, the current box design seems to be better than the Guralp Spec sheet noise by ~a factor of 10.  It would be nice if that number were more like a factor of 100.  Or at least 30.  So some work needs to be done to find a lower-noise op amp for the voltage buffer (the first op amp in the circuit).

Next steps:

Since Larisa is now starting her SURF project with Tara and Mingyuan, I'll look into improving the design of this box by a factor of 3 or 10. 

Then I'll need to make a mock-up of it, and test it out. 

If successful, then I'll draw it up in Altium and have it made.  Recall that there should be 2 outputs per seismometer channel, one with high gain, one with low gain.  Then 3 seismometer channels per seismometer (X, Y, Z), and perhaps multiple seismometer inputs per box.  So lots and lots of stuff all in the same box.  It's going to be pretty cool.

  4809   Mon Jun 13 15:33:55 2011 Jamie, JoeUpdateCDSDolphin fiber between 1Y3 and 1X4 appears to be dead

The fiber that connects the Dolphin card in the c1lsc machine (in the 1Y3 rack) to the Dolphin switch in the 1X4 rack appears to have died spontaneously this morning.  This was indicated by loss of Dolphin communication between c1lsc and c1sus.

We definitively tracked it down to the fiber by moving the c1lsc machine over to 1X4 and testing the connection with a short cable.  This worked fine.  Moving it back to using the fiber again failed.

Unfortunately, we have no replaced Dolphin fiber.  As a work around, we are stealing  a long computer->IO chassis cable from Downs and moving the c1lsc machine to 1X4.

This is will be a permanent reconfiguration.  The original plan was to have the c1lsc machine also live in 1X4.  The new setup will put the computer farther from the RF electronics, and more closely mimic the configuration at the sites, both of which are good things.

  4808   Mon Jun 13 12:34:21 2011 Jamie, KojiUpdateLSCUpdated LSC model installed

After a couple of hickups, I was able to compile and install Koji's rework of the LSC model.

The main changes are that the model now use an RF_PD library part, and the channel names were tweaked to be more in line with what we expect.

I found a couple of small bugs in the model that were preventing it from compiling.  Those were fixed and it compiled with no further problems.

There was also some rearrangement of signal inputs to the PD_DOF matrix.  The matrix screen was updated to reflect the proper inputs.  However, this also meant that the burt restore scripts for the IFO configurations were setting the wrong elements in the matrix.  I fixed the settings for X and Y arm locking, and updated the burt snapshots using the burt/c1ifoconfigure/C1save{X,Y}arm scripts.  NOTE: burt settings will need to be updated for the MICH, PRM, DRM, and FULL IFO configurations as well.

During the build/install process, Joe and I also found a bug in the feCodeGen that was causing the filter screens to be created with the wrong names.  Joe sent out a patch that will hopefully be merged soon.  Building the model with Joe's patch fixed the screen names, so the screens are currently named correctly.

  4807   Fri Jun 10 20:23:56 2011 Larisa ThorneUpdateElectronicsSeismometer Box Update/graphs

 (continuation of this)

 

The noise graphs relating total noise of the Seismometer circuit (GURALP stuff) to the LIGO seismic noise curve have been completed started.

 

 

I apparently harbor hate towards Matlab (you may have notice I do everything in Mathematica)....I will try to change my ways  DX

  4806   Fri Jun 10 18:49:40 2011 kiwamuUpdateIOOIntensity Noise after the MC

Last night the relative intensity noise (RIN) of the beam after MC was measured.

It looks like the RIN is dominated by the motion of the MC mirrors, possibly the angular motions because we don't have any angular stabilization servos.

Suresh will estimate the contribution from the MC mirrors' angular motion to the RIN in order to see if this plot makes sense.

 

(RIN)

 The spectrum below 30 Hz seems to be dominated by the motion of the MC suspensions because it very resembles the spectra of those.

Above 30 Hz the spectrum becomes somehow flat, which I don't know why at the moment.

A rough estimation of the shot noise gave me a level of 10^{-9} RIN/sqrtHz, which is way below the measured spectrum.

RIN_afterMC.png

 

(Setup)

 All of the suspended mirrors were intentionally misaligned except for the MC mirrors and PRM to avoid interference from the other optics.

In this setup it allows us to measure the intensity noise of the laser which is transmitted from MC.

The beam transmitted from MC is reflected by PRM and finally enters into the REFL11 RFPD.

The DC signal from the RFPD was acquired at C1:LSC-REFL_DC_DQ as the laser intensity.

As well as the RIN measurement I took a spectrum when the beam is blocked by a mechanical shutter on the PSL table.

This data contains the dark noise from REFL11 and circuit noise from the whitening, AA board and ADC. It is drawn in black in the plot.

The cut off at 15 Hz is possibly due to the digital unwhitening (two poles at 15 Hz and two zeros at 150 Hz) filter to correct the analog whitening filter.

ELOG V3.1.3-